# TinyR1-32B-Preview: Boosting Accuracy with Branch-Merge Distillation

### v1

@@ 2.2 The Merge Phase
In the Merge phase, we use Arcee Fusion Goddard et al. (2024) to merge models from different domains. This technique selectively integrates meaningful parameter updates from a teacher model $T$ into a student model $S$, rather than simply averaging parameters. It proceeds in three stages:
@@
    Importance Scoring: Compute the importance of each parameter as follows: First, compute the parameter difference $T-S$ and the KL divergence $\operatorname*{KL}(T,S)$. Next, obtain the Importance Score (IS) by multiplying them:
    $\operatorname*{IS}=\operatorname*{KL}(T,S)\cdot(T-S).$
@@
    Dynamic Selection: Determine a threshold $\operatorname*{THR}$ of mask via the median $\operatorname*{Med}$ and interquartile range $\operatorname*{IR}$ of the importance scores:
    $\operatorname*{THR}=\operatorname*{Med}+\theta\cdot\operatorname*{IR},$
    where $\theta$ is a balancing coefficient. Since THR is derived from IS statistics, it dynamically adapts to model parameters, ensuring an optimal update ratio. MergeKit’s developers determined through extensive experiments that the optimal value for $\theta$ is 1.5.
@@
    Selective Integration: Apply a merging mask $\operatorname*{Mask}$ to integrate only parameters exceeding the threshold $\operatorname*{THR}$, outputting the merged model $M$:
    $M=S+\operatorname*{Mask}(T-S).\quad\operatorname*{Mask}=\mathbb{I}_{\operatorname*{IS}>\operatorname*{THR}}.$
    Only teacher parameters with an importance score above the threshold are retained; otherwise, student model parameters are kept.

### v2

@@ 2.2 The Merge Phase
In the Merge phase, we use Arcee Fusion Goddard et al. (2024) to merge models from different domains.
Two different models are utilized during the Merge Phase. We refer to them as the Left Model and the Right Model, with their parameters denoted by $\theta_{L}\in\mathbb{R}^{M}$ and $\theta_{R}\in\mathbb{R}^{M}$, where $M$ is the number of model parameters, respectively.
The merge process is divided into three steps:
@@
    Computing Importance Score
    To measure the impact of each parameter on model performance, an Importance Score $\mathcal{S}_{IS}\in\mathbb{R}^{M}$ is defined, and each $\mathcal{S}_{IS}^{(i)}$ represents the importance of the $i$-th parameter.
    The Importance Score Matrix $S_{IS}$ is computed based on the Kullback-Leibler (KL) divergence between the softmax-transformed samples drawn from the parameter distributions of the two models. Specifically, we first sample two sets of parameters $X$ and $Y$ from the distributions of the two models:
    $\begin{split}X=\{x_{1},x_{2},\dots,x_{N}\},\quad x_{i}\sim p(\theta_{L}),\\ Y=\{y_{1},y_{2},\dots,y_{N}\},\quad y_{i}\sim p(\theta_{R}).\end{split}$ (1)
    where $p(\theta_{L})$ and $p(\theta_{R})$ represent the parameter distributions of the two models. Next, we apply the softmax function to $X$ and $Y$ to convert them into two new distributions named $\tilde{X}$ and $\tilde{Y}$, the definition is shown below:
    $\tilde{x}_{i}=\frac{e^{x_{i}}}{\sum_{j=1}^{N}e^{x_{j}}}+\epsilon,\tilde{y}_{i}=\frac{e^{y_{i}}}{\sum_{j=1}^{N}e^{y_{j}}}+\epsilon,$
    where $\epsilon=1e-8$ is a hyperparameter. The $\mathcal{S}_{IS}$ is then calculated as:
    $\mathcal{S}_{IS}=D_{KL}(\tilde{X}\parallel\tilde{Y})\cdot(\theta_{L}-\theta_{R}).$ (2)
@@
    Calculating Dynamic Selection:
    We sorted $S_{IS}$ and define the following variables:
    • The median $\mathcal{Q}_{Med}$: The middle value of the sorted $S_{IS}$.
    • The lower quartile $Q_{\frac{3}{4}}$: The median of the lower half of the sorted $S_{IS}$.
    • The upper quartile $Q_{\frac{1}{4}}$: The median of the upper half of the sorted $S_{IS}$.
    • The interquartile range $\mathcal{Q}_{IR}$: Defined as $Q_{\frac{3}{4}}-Q_{\frac{1}{4}}$, where $Q_{\frac{3}{4}}\geqslant Q_{\frac{1}{4}}$, ensuring $\mathcal{Q}_{IR}\geqslant 0$.
    $\mathcal{S}_{THR}\in\mathbb{R}$ is defined via the median $\mathcal{Q}_{Med}$ and interquartile range $\mathcal{Q}_{IR}$ :
    $\mathcal{S}_{THR}=\mathcal{Q}_{Med}+\lambda\cdot\mathcal{Q}_{IR},$ (3)
    where $\lambda$ is a hyperparameter defined to balance coefficient.
    Since $\mathcal{S}_{THR}$ is derived from IS statistics, it dynamically adapts to model parameters, ensuring an optimal update ratio.
@@
    Selective Integration:
    The Merged Model is defined as $\theta_{M}\in\mathbb{R}^{M}$. We applied the following rule to merge $\theta_{B}$ and $\theta_{P}$:
    $\theta_{M}^{i}=\theta_{L}^{i}+(\theta_{R}^{i}-\theta_{L}^{i})\cdot\lceil\max(0,\mathcal{S}_{IS}^{i}-\mathcal{S}_{THR})\rceil.$ (4)
    Only Right Model’s parameters with an importance score above the threshold are retained; otherwise, Left Model’s parameters are kept.